How do you put this into symbols and numbers?
Four times the greater of two integers is one more than five times the lesser number. ?

Answer by Edwin McCravy(20086)


Let the greater number be x

Let the lesser number be y

Four times the greater is 1 more than five times the lesser.

   4 times        x    is 1 more than   5  times the lesser.

            4x          = 1 more than   5y                

            4x          =               5y + 1 

                     4x = 5y + 1 

-----------------------------------------------------------
Or

             4(greater) = 5(lesser) + 1

                     4x = 5y + 1

Here are 10 different solutions to that equation:

Greater number = x = 4, Lesser number = y = 3 since 4(4) = 5(3) + 1.
Greater number = x = 9, Lesser number = y = 7 since 4(9) = 5(7) + 1.
Greater number = x = 14, Lesser number = y = 11 since 4(14) = 5(11) + 1.
Greater number = x = 19, Lesser number = y = 15 since 4(19) = 5(15) + 1.
Greater number = x = 24, Lesser number = y = 19 since 4(24) = 5(19) + 1.
Greater number = x = 29, Lesser number = y = 23 since 4(29) = 5(23) + 1.
Greater number = x = 34, Lesser number = y = 27 since 4(34) = 5(27) + 1.
Greater number = x = 39, Lesser number = y = 31 since 4(39) = 5(31) + 1.
Greater number = x = 44, Lesser number = y = 35 since 4(44) = 5(35) + 1.
Greater number = x = 49, Lesser number = y = 39 since 4(49) = 5(39) + 1.

Edwin